OpenStudy (oneowone):

When two bar magnets attract, Do they generate a force on each other equally or just one magnet applying a force of attraction to another magnet? I feel there are multiple forces depending on the amount of magnets to add up to a net total force. What do you all think?

OpenStudy (oneowone):

@radar @Algebraic! @amistre64 @Carl_Pham

OpenStudy (radar):

As you know the force can act in two ways, attraction or repulsion depending on the orientation of the poles of the magnets. I think the total force is the result of both magnets and their respective strength. But this is just my own opinion.

OpenStudy (amistre64):

isnt the set up similar to Gravity?\[F=G\frac{Mm}{d^2}\]but with magnet terms?

OpenStudy (oneowone):

@amistre64 I'm not really sure, I was just wondering do they both equally apply their strength on each other to equal a net total force? @radar Yea I know, in that case if we have two opposite poles that would mean that they attract, I was wondering that force of attraction is applied from BOTH magnets or just one magnet acting on the other? Logically I think both magnets apply that force to add up to a net force because magnets can't just be forced ON by they also force with :P If that makes any sense at all...

OpenStudy (radar):

I don't believe you can separate this, it takes two poles, and they contribute in accordance to their strength. I believe amistre64 shows it is the product of their strength.

OpenStudy (oneowone):

@radar @amistre64 I would agree if I would've understood the law i the beginning... So we can agree that YES the two magnets equally apply their attraction/repulsion force on each other?

OpenStudy (vincent-lyon.fr):

Of course they do! Newton's third law does apply to ALL interactions, including the electromagnetic one.

OpenStudy (vincent-lyon.fr):

When magnets are in line, interaction force is : \(F_{1/2}=\Large \frac {\mu_o}{4\pi}\frac {6M_1M_2}{z^4}\) where \(M_1\) and \(M_2\) are the respective magnetic moments of the magnets. You notice that \(M_1\) and \(M_2\) are interchangeable, confirming Newton's third law, as \(F_{1/2}=F_{2/1}\)

OpenStudy (anonymous):

yes.. i agree with above.. but i wanna extend this question.. what about gravity as defined by einstien.. !! he said gravity is pseudo force.. experienced due to curvature of space time.. well in that cause... i really don't understand how sun experiences the same force due to earth as earth does due to sun??

OpenStudy (anonymous):